### 15 Mar 2026 v1

@@ 7 Case Studies: Poet Profiles
The case studies convert the comparative field into literary argument. Ten poets belong to the balanced inferential cohort and therefore support controlled or within-meter interpretation. Three further anchors, Ferdowsi, Shahriar, and Khayyam, enter descriptively from the wider poet atlas because any long historical account of Persian poetic sound would be distorted without epic, modern lyric, and rubai as reference points. The aim is not to turn every poet into a miniature coefficient table, but to read phonetic profiles alongside genre, performance context, and literary-historical reputation. The broader fingerprint atlas for the full comparative field appears in Appendix A.
@@
Figure 7 provides the visual entry point for the case analyses. Because every poet is plotted on identical axes, the central trade-offs can be read before the prose narrows to individual authors. The figure also clarifies why a single hard-soft vocabulary is inadequate: Hafez and Shah Nematollah Vali are both open, but not in the same way; Ferdowsi and Saadi are both firm, but for different literary functions; Safi Alishah and Khaqani both intensify contour, but not for the same rhetorical ends.
@@
Table 5 summarizes only the ten balanced case-study poets for whom within-meter estimation is adequately supported. Ferdowsi, Shahriar, and Khayyam remain in the section as historically necessary descriptive anchors, but they are not folded into unsupported inferential claims. The distinction matters because the argument of the article depends on keeping literary centrality and evidential support in view together.
Figure 8: Within-meter comparison for the supported case-study subset. Columns correspond to the retained classical meters mutaqarib, hazaj, ramal, mujtass, and muzari’, and rows to focal poets with sufficient mesra support in at least one cell. The three aligned heatmaps report standardized within-meter effects for hardness, sonority, and sibilance; red cells indicate positive deviations, blue cells negative deviations, and gray cells unsupported poet-meter combinations. The figure makes visible Saadi’s portable hardness, Bidel’s multiaxial contrast across several meters, Shah Nematollah Vali’s strong ramal reversal toward low hardness and high sonority, and the more localized contour signatures of Hafez and Safi Alishah.
@@ 8.1 Style as Configuration Rather Than Scalar Rank
One of the main interpretive gains of the article is the shift from scalar ranking to profile configuration. Terms such as smoothness, severity, luminosity, density, or brilliance are common in Persian literary criticism, but they compress several distinct phonetic routes. A poet may sound forceful through hardness, through sibilant contour, through cluster pressure, or through a combination of these. A poet may sound open through sonority, through vowel spread, through reduced compaction, or through all three.
The fingerprints make those routes visible. Saadi’s authority is hardness-led rather than strongly sibilant. Hafez’s lyric memorability depends on openness plus contour. Shah Nematollah Vali and Jahan Malek Khatun share sonorous breadth, but not the same hardness reversal or entropy. Bidel’s difficulty arises less from blunt roughness than from internally varied pressure. Ferdowsi shows that epic authority can remain forceful without bright fricative emphasis. The value of measurement here is not to replace critical vocabulary, but to decompose it into more exact comparative terms.
@@ 8.4 Implications for Digital Humanities
For digital humanities, the broader lesson is methodological. A strong humanities-computation article should not oppose reproducibility to interpretation. It should make representation choices explicit, control design historically intelligible, and figures readable enough for scrutiny. Quantitative work is most persuasive when it identifies which differences are formally structured, historically meaningful, and worth renewed critical debate.
In that sense, the contribution of this article is not only substantive but programmatic. It suggests a way of studying poetic sound that is cumulative rather than reductive: corpus-scale enough to reveal structure, controlled enough to support comparison, and interpretive enough to remain answerable to literary history.
@@ 9 Limitations
The strongest claims of this article are deliberately tied to a restricted inferential domain. Controlled models are estimated only on the five most frequent named classical meters and on poet-meter cells with at least 2,000 mesras, so low-support poets, rare forms, and some historically important corpora fall outside the main comparison set. Ferdowsi, Shahriar, and Khayyam illustrate this boundary directly: they remain indispensable for comparison, but they should not be absorbed into the same within-meter inferential frame as the high-support cohort. The resulting effect sizes are meaningful yet moderate, and they should be read as structured stylistic tendencies rather than as exhaustive summaries of poetic value.
The analysis is also mediated by representation and corpus history. Persian orthography omits many short vowels, so the grapheme-to-phoneme layer remains a disciplined reconstruction rather than a record of historical recitation; meter, form, and line length are controlled, but finer performative and regional conditions are not. Attribution uncertainty, genre concentration, and canon formation further shape which poets survive in volumes large enough for stable estimation. The framework is therefore best understood as a confound-aware comparative instrument that invites future genre-stratified, attribution-sensitive, and performance-aware extension.
@@ 10 Conclusion
This article has argued that phonetic texture in Persian poetry can be studied as a serious digital-humanities object when representation, confound control, and literary interpretation are kept in view together. The central empirical finding is not that poets occupy acoustically isolated worlds, but that measurable phonetic differentiation persists within shared formal institutions. Meter and form matter strongly; poet identity still matters after they are controlled. The result is a historically regulated field of variation rather than either formal determinism or authorial essence.
The study therefore contributes on four linked fronts already outlined in the introduction: a corpus-scale phonetic analysis of Persian poetic language, a confound-aware framework for stylistic comparison, a fingerprint-based account of poet-level configuration, and a reproducible digital-humanities workflow that keeps interpretive stakes visible. Empirically, the article shows that Persian poetic sound can be compared systematically without collapsing it into one index, that poet profiles are organized by trade-offs among openness, hardness, contour, entropy, and compaction within recognizable aruz repertoires, and that century-level change is better described as redistribution than rupture.
@@
For digital humanities, the broader implication is methodological. A rigorous computational poetics of Persian literature should not oppose measurement to interpretation. It should use measurement to discipline interpretation, and interpretation to keep measurement historically meaningful. Under that standard, phonetic analysis becomes more than a technical supplement to stylometry: it becomes a way of hearing the corpus historically.
The restriction to five meters is therefore substantive rather than merely procedural. The retained bahrsmutaqarib, hazaj, ramal, mujtass, and muzari’—are historically documented classical meters, not ad hoc metrical families, and together they account for 77.6% of mesras with usable meter/form metadata and the entirety of the final controlled cohort. That coverage makes them the appropriate basis for corpus-wide comparison in this article: they dominate the metrically annotated archive, span epic, romance, lyric, mystical, rhetorical, and revivalist uses, and support generalization without collapsing rarer meters into spurious aggregate categories.
@@ Appendix A Poet Fingerprint Atlas
This appendix presents the broader fingerprint atlas used to situate the main case studies against the wider comparative field. Each sheet preserves a shared radial scale across hardness, sonority, sibilance, vowel ratio, entropy, and consonant-cluster ratio, while grouping between eight and twelve poets per page to keep the atlas readable at journal scale. Read the pages comparatively: the point is not to isolate single poets, but to see how epic, lyric, courtly, mystical, revival, and modern corpora redistribute the same phonetic resources.
@@ Appendix B Meter-Stratified Coefficient Tables
These tables report within-meter standardized poet effects for the three focal metrics. They are cited in the main text whenever meter portability is discussed and are included here so that readers can inspect the evidential base behind the summary statements on hardness, sonority, and sibilance across mutaqarib, hazaj, ramal, mujtass, and muzari’.

### 01 Apr 2026 v2

@@ 7 Case Studies: Poet Profiles
The case studies place the comparative field in literary-historical context. Ten poets belong to the balanced inferential cohort and therefore support controlled or within-meter interpretation. Three further anchors, Ferdowsi, Shahriar, and Khayyam, enter descriptively from the wider poet atlas because any long historical account of Persian poetic sound would be distorted without epic, modern lyric, and rubai as reference points. Throughout the section, phonetic profiles are read alongside genre, performance context, and literary-historical reputation. The broader fingerprint atlas for the full comparative field appears in Appendix A.
@@
Figure 7 sets the case analyses on a common comparative scale. Because every poet is plotted on identical axes, the central trade-offs emerge before the discussion narrows to individual authors. The figure also clarifies why a single hard-soft vocabulary is inadequate: Hafez and Shah Nematollah Vali are both open, but not in the same way; Ferdowsi and Saadi are both firm, but for different literary functions; Safi Alishah and Khaqani both intensify contour, but not for the same rhetorical ends.
@@
Table 5 summarizes only the ten balanced case-study poets for whom within-meter estimation is adequately supported. Ferdowsi, Shahriar, and Khayyam remain in the section as historically necessary descriptive anchors, but they are not folded into unsupported inferential claims. The distinction matters because the discussion keeps literary centrality and evidential support in view together.
Figure 8: Within-meter comparison for the supported case-study subset. Columns correspond to the retained classical meters mutaqarib, hazaj, ramal, mujtass, and muzari’, and rows to focal poets with sufficient mesra support in at least one cell. The three aligned heatmaps report standardized within-meter effects for hardness, sonority, and sibilance; red cells indicate positive deviations, blue cells negative deviations, and gray cells unsupported poet-meter combinations. The figure highlights Saadi’s portable hardness, Bidel’s multiaxial contrast across several meters, Shah Nematollah Vali’s strong ramal reversal toward low hardness and high sonority, and the more localized contour patterns of Hafez and Safi Alishah.
@@ 8.1 Style as Configuration Rather Than Scalar Rank
The analysis shifts attention from scalar ranking to profile configuration. Terms such as smoothness, severity, luminosity, density, or brilliance are common in Persian literary criticism, but they compress several distinct phonetic routes. A poet may sound forceful through hardness, through sibilant contour, through cluster pressure, or through a combination of these. A poet may sound open through sonority, through vowel spread, through reduced compaction, or through all three.
The phonetic fingerprints make those routes easier to distinguish. Saadi’s authority is hardness-led rather than strongly sibilant. Hafez’s lyric memorability depends on openness plus contour. Shah Nematollah Vali and Jahan Malek Khatun share sonorous breadth, but not the same hardness reversal or entropy. Bidel’s difficulty arises less from blunt roughness than from internally varied pressure. Ferdowsi shows that epic authority can remain forceful without bright fricative emphasis. The value of measurement here is not to replace critical vocabulary, but to decompose it into more exact comparative terms.
@@ 8.4 Implications for Digital Humanities
For digital humanities, the broader lesson is methodological. Reproducibility and interpretation do not have to compete. Representation choices should be explicit, the control design should remain historically intelligible, and figures should support close scrutiny. Quantitative work is most persuasive when it identifies which differences are formally structured, historically meaningful, and worth renewed critical debate.
At corpus scale, this approach studies poetic sound without flattening literary history. It is cumulative rather than reductive: broad enough to reveal structure, controlled enough to support comparison, and interpretive enough to remain answerable to literary history.
@@ 9 Limitations
The strongest claims are tied to a restricted inferential domain. Controlled models are estimated only on the five most frequent named classical meters and on poet-meter cells with at least 2,000 mesras, so low-support poets, rare forms, and some historically important corpora fall outside the main comparison set. Ferdowsi, Shahriar, and Khayyam illustrate this boundary directly: they remain indispensable for comparison, but they should not be absorbed into the same within-meter inferential frame as the high-support cohort. The resulting effect sizes are meaningful yet moderate, and they should be read as structured stylistic tendencies rather than as exhaustive summaries of poetic value.
The analysis is also mediated by representation and corpus history. Persian orthography omits many short vowels, so the grapheme-to-phoneme layer remains a disciplined reconstruction rather than a record of historical recitation; meter, form, and line length are controlled, but finer performative and regional conditions are not. Attribution uncertainty, genre concentration, and canon formation further shape which poets survive in volumes large enough for stable estimation. The framework remains a confound-aware comparative instrument that invites future genre-stratified, attribution-sensitive, and performance-aware extension.
@@ 10 Conclusion
This study shows that phonetic texture in Persian poetry can be studied systematically when representation, confound control, and literary interpretation are kept in view together. The central empirical finding is not that poets occupy acoustically isolated worlds, but that measurable phonetic differentiation persists within shared formal institutions. Meter and form matter strongly, yet poet identity still matters after they are controlled. The result is a historically regulated field of variation rather than either formal determinism or authorial essence.
Taken together, the analysis provides a corpus-scale phonetic account of Persian poetic language, a confound-aware framework for stylistic comparison, and a fingerprint-based view of poet-level configuration within recognizable aruz repertoires. The workflow remains reproducible and interpretively accountable. Empirically, Persian poetic sound can be compared systematically without collapsing it into one index, and century-level change is better described as redistribution than rupture.
@@
For digital humanities, the broader implication is methodological. Measurement and interpretation are most useful when they discipline one another. Under that standard, phonetic analysis becomes more than a technical supplement to stylometry: it becomes a way of hearing the corpus historically.
The restriction to five meters is therefore substantive rather than merely procedural. The retained bahrsmutaqarib, hazaj, ramal, mujtass, and muzari’—are historically documented classical meters, not ad hoc metrical families, and together they account for 77.6% of mesras with usable meter/form metadata and the entirety of the final controlled cohort. That coverage makes them the appropriate basis for corpus-wide comparison here: they dominate the metrically annotated archive, span epic, romance, lyric, mystical, rhetorical, and revivalist uses, and support generalization without collapsing rarer meters into spurious aggregate categories.
@@ Appendix A Poet Fingerprint Atlas
This appendix presents the broader fingerprint atlas used to situate the main case studies against the wider comparative field. Each sheet preserves a shared radial scale across hardness, sonority, sibilance, vowel ratio, entropy, and consonant-cluster ratio, while grouping between eight and twelve poets per page. Taken together, the pages show how epic, lyric, courtly, mystical, revival, and modern corpora redistribute the same phonetic resources across the wider field rather than isolating single poets as self-contained cases.
@@ Appendix B Meter-Stratified Coefficient Tables
These tables report within-meter standardized poet effects for the three focal metrics. They provide the evidential base for the summary statements on hardness, sonority, and sibilance across mutaqarib, hazaj, ramal, mujtass, and muzari’.